Trying to understand inverse demand. If my demand curve is Qx = 400 + 0.0021(I)+ 8(Z) - 5(Py)- 20(Px) where I=10,000; Z= 150; Py=50; Px = 50. Resulting in Qx = 370. How to solve for the inverse?
Answer #1

First note that the demand for X is a function of its price Px and other factors. So use the values of all other factors in the demand function except PX.

QX = 400 + 0.0021*10000 + 8*150 - 5*50 - 20PX

QX = 1371 - 20PX

Now the inverse demand function is

20PX = 1371 - QX

PX = 1371/20 - QX/20

PX = 68.55 - 0.05QX